Let us assume that Fe(OH)3(s) is completely insoluble, which signifies that the precipitation reaction with NaOH(aq) (presented in the transition) would go to completion.

Fe3+(aq)+3NaOH(aq) ? Fe(OH)3(s)+3Na+(aq)

If you had a 0.450 L solution containing 0.0200 M of Fe3+(aq), and you wished to add enough 1.33 M  NaOH(aq) to precipitate all of the metal, what is the minimum amount of the NaOH(aq)solution you would need to add? Assume that the NaOH(aq) solution is the only source of OH?(aq) for the precipitation.

Express the volume to three significant figures and include the appropriate units

minimal volume of NaOH(aq) solution =
Answer #1

Fe3+(aq) + 3NaOH(aq) = Fe(OH)3(s) + 3Na+(aq)

(0.450L) * ( 0.0200 M Fe{3+} * 3mol NaOH / 1 mol Fe{3+}) / ( 1.33 mol / L NaOH ) = 0.0203 L = 20.3 mL NaOH

Three significant figures